# Windowed MAPF with Completeness Guarantees

### v1

@@ Abstract
Traditional multi-agent path finding (MAPF) methods try to compute entire start-goal paths which are collision free. However, computing an entire path can take too long for MAPF systems where agents need to replan fast. Methods that address this typically employ a “windowed” approach and only try to find collision free paths for a small windowed timestep horizon. This adaptation comes at the cost of incompleteness; all current windowed approaches can become stuck in deadlock or livelock. Our main contribution is to introduce our framework, WinC-MAPF, for Windowed MAPF that enables completeness. Our framework uses heuristic update insights from single-agent real-time heuristic search algorithms as well as agent independence ideas from MAPF algorithms. We also develop Single-Step CBS (SS-CBS), an instantiation of this framework using a novel modification to CBS. We show how SS-CBS, which only plans a single step and updates heuristics, can effectively solve tough scenarios where existing windowed approaches fail.
@@ 1 Introduction
A core problem for multi-agent systems is to figure out how agents should move from their current location to their goal location. Without careful consideration, agents can collide, get stuck in deadlock, or take inefficient paths which take longer to traverse. This Multi-Agent Path Finding (MAPF) problem is particularly tough in congestion or when the number of agents becomes very large (e.g. 100s).
Initially, most MAPF methods attempted to find entire paths for each agent from their start to their goal. These solvers could take tens of seconds to even minutes to find a collision-free solution. In practice, this can require agents to idle during the planning phase and only execute once the solution is found. Real-world practitioners and systems aim to avoid having agents idle unnecessarily during long planning phases. Therefore instead of finding an entire collision-free path, newer methods took existing full-horizon MAPF solvers and modified them to only reason about collisions within a limited horizon.
Concretely, these methods typically define a fixed time window $W$ and plan paths for each agent to the goal such that the first $W$ timesteps account for inter-agent coordination and avoid collisions. This window $W$ is typically much shorter than the entire solution path, e.g. $W=5$ is common when the entire solution path spans 50 to 500 timesteps. As a result, windowed methods are significantly faster than those that compute the entire path.
A key issue with these windowed approaches is that their myopic planning results in deadlock or livelock if their window is too small. Table 1 shows examples where windowed MAPF solvers fail if their window is too small in congestion. More broadly, all existing windowed MAPF solvers regardless of window size lack theoretical completeness and several windowed works have explicitly cited deadlock as a key issue in their experiments (Li et al. 2020; Okumura et al. 2022; Jiang et al. 2024).
@@
Table 1: Success rates of different windowed optimal approaches (i.e. ECBS and EECBS run with suboptimality=1) on small congested scenarios (with varying numbers of agents) across 20 seeds and a 1 minute timeout. Windowed approaches like ECBS fail due to deadlock or timeout. Even full horizon EECBS with all optimizations fails due to timeouts. SS-CBS, with a single-step planning horizon, is able to solve these scenarios using our WinC-MAPF framework which maintains completeness guarantees.
Our first main contribution is the introduction of the Windowed Complete MAPF framework, WinC-MAPF, designed to create Windowed MAPF solvers that guarantee completeness. WinC-MAPF is a general framework that leverages concepts from single-agent heuristic search and the semi-independent structure of MAPF. First, we view Windowed MAPF in its joint-configuration and show how we can apply real-time heuristic updates on the joint-configuration to enable completeness. However, due to the large joint state space in a MAPF problem, naive heuristic updates are infeasible. Thus, second, we leverage the semi-independent structure of MAPF problems to focus heuristic updates on different groups of coupled agents simultaneously, resulting in efficient performance. An important module in the WinC-MAPF framework is an Action Generator (AG) that computes the next set of actions the agents need to execute. To guarantee completeness, the AG must accurately identify coupled agents and use the updated heuristic values of different states.
To this extent, our second main contribution is developing Single-Step CBS (SS-CBS), a CBS-based AG that follows the WinC-MAPF framework and plans only for a single timestep ($W=1$). We first how naively integrating heuristic updates in CBS can fail. Thus, SS-CBS introduces a novel “heuristic conflict” and constraint to address this issue. We empirically demonstrate how SS-CBS, with single step planning, outperforms windowed ECBS with significantly larger windows across both small and large instances.
@@ 2.1 Problem Formulation
Multi-Agent Path Finding (MAPF) is the problem of finding collision-free paths for a group of $N$ agents ${i=1,\dots,N}$, that takes each agent from its start location $s_{i}^{\text{start}}$ to its goal location $s_{i}^{\text{goal}}$. In traditional 2D MAPF, the environment is discretized into grid cells, and time is broken down into discrete timesteps. Agents are allowed to move in any cardinal direction or wait in the same cell. A valid solution is a set of agent paths $\Pi=\{\pi_{1},...,\pi_{N}\}$ where $\pi_{i}[0]=s_{i}^{start}$, $\pi_{i}[T_{i}]=s_{i}^{goal}$ where $T_{i}$ is the maximum timestep of the path for agent $i$. Critically, agents must avoid vertex collisions (when $\pi_{i}[t]=\pi_{j\neq i}[t]$) and edge collisions (when $\pi_{i}[t]=\pi_{j}[t+1]\wedge\pi_{i}[t+1]=\pi_{j}[t]$) for all $t$ timesteps. The typical objective in optimal MAPF is to find a solution $\Pi$ that minimizes $|\Pi|=\sum_{i=1}^{N}|\pi_{i}|=\sum_{i=1}^{N}\sum_{t=0}^{T_{i}-1}c(s_{i}^{t},s_{i}^{t+1})$.
@@
Planning a set of full horizon collision-free paths can take on the order of 10s of seconds, which is too slow for some applications. For instance, a recent MAPF competition, League of Robot Runners (Chan et al. 2024), required planning for hundreds of agents within 1 second. Therefore, instead of planning full horizon collision-free paths, several methods (discussed in the next section) only resolve collisions for a fixed windowed horizon $W$. As a result, instead of planning full-horizon collision-free paths, several methods (discussed in the next section) focus on resolving collisions only within a fixed windowed horizon $W$. Concretely, instead of reasoning about collisions for all $t$, these methods only reason about collisions for $t\leq W$. Thus after the first $W$ timesteps, the remaining path $\pi_{i}[W]...\pi_{i}[T_{i}]$ will simply be the agent’s optimal path to the goal as it does not need to avoid collisions with other agents.
Mathematically then, the cost of $\pi_{i}$ is $\sum_{t=0}^{W-1}c(s_{i}^{t},s_{i}^{t+1})+c^{*}(s_{i}^{W},s_{i}^{goal}).$ Additionally, all performant 2D MAPF methods compute a backward dijkstra’s for each agent where $h^{*}_{i}(s)=c^{*}(s,s_{i}^{goal})$. Thus instead of planning each $\pi_{i}^{0:T_{i}}$, we can equivalently plan just the windowed horizon $\pi_{i}^{0:W}$ and minimize the same $\Pi$ objective which is now equal to $\sum_{i}^{N}(\sum_{t=0}^{W-1}c(s_{i}^{t},s_{i}^{t+1})+h^{*}_{i}(s_{i}^{W}))$.
@@ 2.2 MAPF Methods
There exist many different types of heuristic search solvers for MAPF. One old approach is Prioritized Planning (Erdmann and Lozano-Perez 1987) which assigns priorities to agents and plans them sequentially with later agents avoiding earlier agents. PIBT (Okumura et al. 2022) is a recent popular method that allows agents to “inherit” other agents’ priorities. Conflict Based Search (Sharon et al. 2015) is another popular method that decoupled the planning problem into two stages. A high-level search resolves conflicts between agents by applying constraints while a low-level search finds plans for individual agents that satisfy constraints. There are many extensions to CBS that improve the searches as well as the applied constraints (Barer et al. 2014; Li, Ruml, and Koenig 2021; Li et al. 2021).
When faced with shorter planning times, methods typically simplify the planning problem to just find partial collision-free paths. Windowed Hierarchical Cooperative A* (Silver 2005) is a windowed variant of Hierarchical Cooperative A* which is essentially a prioritized planner using a backward Dijkstra’s heuristic. Silver notes that this is not complete due to their use of priorities. Rolling Horizon Conflict Resolution (RHCR) applies a rolling horizon for lifelong MAPF planning and replans paths at repeated intervals (Li et al. 2020). RHCR faces deadlock and attempts to combat it by increasing the planning window but still notes that their method is incomplete. Bounded Multi-Agent A* (Sigurdson et al. 2018) proposes that each agent runs its own limited horizon real-time planner considering other agents as dynamic obstacles. However, the method acknowledges that deadlock can become a problem when agents need to coordinate with one another.
Planning and Improving while Executing (Zhang et al. 2024) is a recent work that attempts to quickly generate an initial full plan using LaCAM (Okumura 2022) and then refines it during execution using LNS2 (Li et al. 2022). However, if a complete plan cannot be found, the method resorts to using the best partial path available, making it incomplete in such situations. The winning submission (Jiang et al. 2024) to the Robot Runners competition, due to the tight planning time constraint, leveraged windowed planning of PIBT with LNS2. They explicitly note deadlock in congestion is a significant challenge.
@@ 2.3 Real-Time Single Agent Search
We leverage ideas from “Real-Time” search, a single-agent heuristic search problem where due to limited time constraints the agent is required to iteratively plan and execute partial paths. Despite repeatedly planning partial paths, Real-Time search methods are designed to maintain completeness. The main innovation in single-agent real-time search literature is that the agent updates (increases) the heuristic value of encountered states. This prevents deadlock/livelock as states that are repeatedly visited have larger and larger heuristic values which encourages the search to explore other areas. A large variety of real-time algorithms such as LRTA* (Korf 1990), RTAA* (Koenig and Likhachev 2006), and LSS-LRTA* (Koenig and Sun 2009) propose to update the heuristic in different ways.
The core idea is that given a current state $s$ and a partial plan leading to a new state $s^{\prime}$, we update the heuristic value as $h(s)\leftarrow c(s,s^{\prime})+h(s^{\prime})$. Thus when stuck in a local minima, the agent repeatedly visits states in the local minima and updates their heuristic values until they become too high and cause the agent to expand states outside the local minima. Most single-agent heuristic search algorithms rely on optimal planners to pick the next state to move to.
@@ 3 Windowed-MAPF with Guarantees
This section describes our Windowed Complete MAPF framework, WinC-MAPF, for creating windowed MAPF solvers that guarantee completeness. We leverage two key insights. Our first insight is that we can apply single-agent real-time update ideas to MAPF planning if we interpret the MAPF problem as a single-agent problem in the combined joint state space. This allows us to update the heuristics of previously seen states enabling completeness. However, just doing this is ineffective due to the large state space. To this extent, our second insight is that we can leverage MAPF’s agent semi-independence to intelligently update the heuristic value of multiple states, allowing the search to fill in heuristic depressions quickly and exit local minima faster.
@@
Figure 1: The iterative planning, execution, and grouped heuristic updates/penalties for Windowed Complete MAPF (WinC-MAPF).
3.1 Planning in Joint-State Space
@@
Our first observation is to leverage existing Real-Time search literature that has been solely explored in single-agent planning as mentioned in Section 2.3. We can directly leverage single-agent Real-Time search literature if we view our multi-agent problem in the joint space.
@@
We formally redefine the windowed MAPF planning problem from this joint-space perspective. Given $N$ agents, we define a joint configuration $\mathcal{C}^{t}=[s_{1}^{t},...s_{N}^{t}]$. At every timestep, we query a high-level “action-generator” to return a sequence of configurations $\Pi^{0:W}=[\mathcal{C}^{0},...\mathcal{C}^{W}]$ which minimizes $|\Pi|=|\Pi^{0:W}|+|\Pi^{W:T_{max}}|=\sum_{t=0}^{W-1}c(\mathcal{C}^{t},\mathcal{C}^{t+1})+h_{BD}(\mathcal{C}^{W})$. We define the joint cost and heuristic intuitively, $c(\mathcal{C},\mathcal{C}^{\prime})=\sum_{i}^{N}c(s_{i},s_{i}^{\prime})$ and $h_{BD}(\mathcal{C})=\sum_{i}^{N}h^{*}_{i}(s_{i})$ (BD = Backward Dijkstra).
To re-iterate, the advantage of this interpretation is that we have converted our windowed MAPF problem into a standard single-agent real-time search problem. Thus given this sequence of configurations, we apply a standard Bellman Update to the heuristic of $\mathcal{C}$ via $h(\mathcal{C})\leftarrow c(\mathcal{C},\mathcal{C}^{H})+h(\mathcal{C}^{H})$. Given a timestep $t$ and configuration $\mathcal{C}^{t}$, we move our agents to $\mathcal{C}^{t+1}$, update our heuristic of $\mathcal{C}^{t}$, and repeat.
@@
$h(\mathcal{C})$ is initially set to $h_{BD}(\mathcal{C})$ and gradually increases as agents visit states. Instead of calling these changes heuristic updates, we use the term “heuristic penalty” to describe how the heuristic for visited states increases when we apply our update equation. This effectively “penalizes” those states and encourages the search to explore other states. A heuristic penalty state is a state $\mathcal{C}$ which has a non-zero increase from the base heuristic value, i.e. $h(\mathcal{C})=h_{BD}(\mathcal{C})+h_{p}(\mathcal{C})$ where the penalty $h_{p}(\mathcal{C})>0$. Our penalty update equation is then $h_{p}(\mathcal{C})=c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})-h_{BD}(\mathcal{C})$.
Now during execution, if the MAPF instance is stuck in deadlock at some location $\mathcal{C}$, we can show that the heuristic $h(\mathcal{C})$ will continually increase. Thus after enough iterations, the action generator which minimizes $\operatorname*{arg\,min}_{\mathcal{C}^{\prime}}c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$ will eventually choose a different state (i.e. the chosen $\mathcal{C}^{\prime}\neq\mathcal{C}$) as $h(\mathcal{C})$ will be too large, and leave the deadlock.
Quickly recapping, we view the windowed MAPF problem as a real-time problem in the joint space, requiring an “action-generator” that reasons about $h(\mathcal{C})=h_{BD}(\mathcal{C})+h_{p}(\mathcal{C})$. However, existing MAPF methods like Prioritized Planning, CBS, and LaCAM do not reason about this $h(\mathcal{C})$ out of the box. We show how this can be done for CBS in Section 4.
3.2 Reasoning about Coupled Agents
The framework we have described so far suffers from an obvious issue; the size of the joint state-space $\mathcal{C}$ is very large. Consequently, escaping local minima, which in our context are agents stuck in deadlock/livelock, can be challenging. The process often requires filling in large heuristic depressions, which can lead to poor performance as the agent may remain stuck in the minimum for extended periods before finding a way out.
Concretely, imagine we have the starting configuration as depicted in Figure 1 (where the goals are indicated with arrows). Note that the blue and green agents on the bottom want to swap locations. The main observation is that the orange agents and the blue/green agents are independent even though they are on the same connected graph. However, iteratively planning and computing heuristic penalties on the joint configurations as described will result in the action generator wanting to avoid the exact joint configurations penalized. Thus, the AG could request the orange agents to move, encountering new configurations without penalties, even though the underlying blue-green deadlock remains. To resolve this deadlock, the current framework requires the AG to search over all agent’s locations even though we intuitively know it should focus on just the blue and green agents.
Our key idea is therefore to apply the heuristic penalty to specific groups of agents instead of on the entire joint-state space. In Figure 1, this means that instead of applying $h$ on the single state $\mathcal{C}$, we can determine the specific groups of agents (e.g. blue and green) and only attribute the heuristic penalty to them. Conceptually, the idea is that if each agent $i$ was able to move from $s$ to their optimal next state $s^{\prime}$, there would be no heuristic penalty as each $h_{i}(s)=c(s,s^{\prime})+h_{i}(s^{\prime})$ due to the heuristic being a perfect single-agent heuristic. Thus, $h_{p}(\mathcal{C})>0$ means there exists at least one agent $i$ with $h_{i}(s)<c(s,s^{\prime})+h_{i}(s^{\prime})$ unable to move to its best location. Additionally, this agent $i$ must be blocked by some other agent $j$. Thus, instead of updating the heuristic of $\mathcal{C}$, we can update $\mathcal{C}_{Gr=\{i,j\}}$ (assuming no other agents are interacting with $i,j$).
Our objective is, given a $\mathcal{C},\mathcal{C}^{\prime}$ transition, to detect these groups of agents and apply heuristic penalties to just the group of agents rather than the entire joint-state space. Concretely, for each agent we determine which other agents it is directly interacting with. Agents that are able to do their best action, and whose best action does not block another agent’s movement, are “independent” to all other agents. On the flip side, agents that cannot do their best action due to another agent are coupled with the impeding agents.

### v2

@@ Abstract
Traditional multi-agent path finding (MAPF) methods try to compute entire collision free start-goal paths, with several algorithms offering completeness guarantees. However, computing partial paths offers significant advantages including faster planning, adaptability to changes, and enabling decentralized planning. Methods that compute partial paths employ a “windowed” approach and only try to find collision free paths for a limited timestep horizon. While this improves flexibility, this adaptation introduces incompleteness; all existing windowed approaches can become stuck in deadlock or livelock. Our main contribution is to introduce our framework, WinC-MAPF, for Windowed MAPF that enables completeness. Our framework leverages heuristic update insights from single-agent real-time heuristic search algorithms and agent independence ideas from MAPF algorithms. We also develop Single-Step Conflict Based Search (SS-CBS), an instantiation of this framework using a novel modification to CBS. We show how SS-CBS, which only plans a single step and updates heuristics, can effectively solve tough scenarios where existing windowed approaches fail.
@@ 1 Introduction
A core problem for multi-agent systems is to figure out how agents should move from their current locations to their goal locations. Without careful consideration, agents can collide, get stuck in deadlock, or take inefficient paths which take longer to traverse. This Multi-Agent Path Finding (MAPF) problem is particularly tough in congestion or when the number of agents becomes very large (e.g. 100s).
Full horizon MAPF methods attempt to find entire paths for each agent from their start to their goal. Many such MAPF methods additionally have theoretical completeness guarantees, i.e., they will find a solution if one exists given enough time and compute. However, planning partial paths has multiple advantages including decreasing planning time, adaptability to changes, and enabling decentralized planning. This is particularly useful in scenarios where planning a full horizon path may be tough (e.g. too many collisions to resolve) and planning a partial path is more feasible.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 1: Success rates of different optimal approaches: windowed CBS (wCBS) and full horizon CBS with all optimizations enabled (CBS+). Horizon denotes window size. We test on small congested scenarios (number of agents written below each map) across 20 seeds and a 1 minute timeout. wCBS and CBS+ fail due to deadlock or timeout while SS-CBS, with a single-step planning horizon, is able to solve these scenarios using our WinC-MAPF framework which maintains completeness guarantees.
Therefore existing windowed methods define a time window $W$ and plan paths for each agent to the goal such that the first $W$ timesteps account for inter-agent coordination and avoids collisions. This window $W$ is typically much shorter than the entire solution path, e.g. $W=5$ is common when the entire solution path spans 50 to 500 timesteps. As a result, windowed methods are significantly faster than those that compute the entire path.
A key issue with windowed approaches is that their myopic planning results in deadlock or livelock. Table 1 shows examples where windowed MAPF solvers fail in congestion if their window is too small. More broadly, all existing windowed MAPF solvers regardless of window size lack theoretical completeness, and several windowed works have explicitly cited deadlock as a key issue in their experiments (Silver 2005; Li et al. 2020; Okumura et al. 2022; Jiang et al. 2024). Interestingly, no previous work has addressed completeness in windowed MAPF solvers.
Our first main contribution is the introduction of the Windowed Complete MAPF framework, WinC-MAPF, designed to create Windowed MAPF solvers that guarantee completeness. WinC-MAPF is a general framework that leverages concepts from single-agent real-time heuristic search and the semi-independent structure of MAPF. First, we view Windowed MAPF in its joint configuration and show how we can apply real-time heuristic updates on the joint configuration to enable completeness. However, due to the large joint configuration space in a MAPF problem, naively applying heuristic updates is impractical. Thus, second, we leverage the semi-independent structure of MAPF problems to focus heuristic updates on disjoint agent groups simultaneously, resulting in efficient performance. An important module in the WinC-MAPF framework is an Action Generator (AG) that computes the next set of actions the agents need to execute. To guarantee completeness, the AG must (1) find the optimal windowed action that incorporates heuristic updates, and (2) identify disjoint agent groups. Our framework and proof encourage future research on the development of windowed AGs that satisfy these properties.
As a starting point, our second main contribution is developing Single-Step Conflict-Based Search (SS-CBS), a CBS-based AG that follows the WinC-MAPF framework and plans only for a single timestep ($W=1$). Conflict-Based Search (Sharon et al. 2015) can be easily modified to be windowed by only considering conflicts within the window, but we show how naively integrating heuristic updates can fail. Thus, SS-CBS introduces a novel “heuristic conflict” and constraint to address this issue. We empirically demonstrate how SS-CBS, with single step planning, outperforms windowed CBS with larger windows across both small and large instances.
@@ 2.1 Problem Formulation
Multi-Agent Path Finding (MAPF) is the problem of finding collision-free paths for a group of $N$ agents ${i=1,\dots,N}$, that takes each agent from its start location $s_{i}^{\text{start}}$ to its goal location $s_{i}^{\text{goal}}$. In traditional 2D MAPF, the environment is discretized into grid cells, and time is broken down into discrete timesteps. Agents are allowed to move in any cardinal direction or wait in the same cell. A valid solution is a set of agent paths $\Pi=\{\pi_{1},...,\pi_{N}\}$ where $\pi_{i}[0]=s_{i}^{start}$, $\pi_{i}[T_{i}]=s_{i}^{goal}$ where $T_{i}$ is the maximum timestep of the path for agent $i$. Critically, agents must avoid vertex collisions (when $\pi_{i}[t]=\pi_{j\neq i}[t]$) and edge collisions (when $\pi_{i}[t]=\pi_{j}[t+1]\wedge\pi_{i}[t+1]=\pi_{j}[t]$) for all timestep $t$. The typical objective in optimal MAPF is to find a solution $\Pi$ that minimizes $|\Pi|=\sum_{i=1}^{N}|\pi_{i}|=\sum_{i=1}^{N}\sum_{t=0}^{T_{i}-1}c(s_{i}^{t},s_{i}^{t+1})$. This work solves standard MAPF which has $c(s_{i}^{t},s_{i}^{t+1})=1$ unless the agent is resting at its goal (where $c(s_{i}^{\text{goal}},s_{i}^{\text{goal}})=0$).
@@
Instead of resolving all collisions, windowed planners iteratively plan a smaller collision-free path and execute, in essence breaking the problem into smaller more feasible chunks. For example, a recent MAPF competition, League of Robot Runners (Chan et al. 2024), required planning for hundreds of agents within 1 second and utilized windowed planning interleaved with execution.
Windowed MAPF methods plan partial paths that only reason about collisions for timesteps $t\leq W$ where $W$ is a hyper-parameter window size. Thus after the first $W$ timesteps, the remaining path $\pi_{i}[W]...\pi_{i}[T_{i}]$ is the agent’s optimal path to the goal (in the absence of other agents) as it does not need to avoid collisions with other agents. Mathematically then, the cost of $\pi_{i}$ is $\sum_{t=0}^{W-1}c(s_{i}^{t},s_{i}^{t+1})+c^{*}(s_{i}^{W},s_{i}^{goal})$ where $c^{*}(s_{i}^{W},s_{i}^{goal})$ is the optimal cost to the goal. We note that all performant 2D MAPF methods compute a backward dijkstra’s for each agent where $h^{*}_{i}(s)=c^{*}(s,s_{i}^{goal})$. Thus instead of fully planning $\pi_{i}^{0:T_{i}}$, we can equivalently plan just the windowed horizon $\pi_{i}^{0:W}$ and minimize $|\Pi|=\sum_{i=1}^{N}(\sum_{t=0}^{W-1}c(s_{i}^{t},s_{i}^{t+1})+h^{*}_{i}(s_{i}^{W}))$.
@@ 2.2 MAPF Methods
There exist many different types of heuristic search solvers for MAPF. One old approach is Prioritized Planning (Erdmann and Lozano-Perez 1987) which assigns priorities to agents and plans them sequentially with later agents avoiding earlier agents. PIBT (Okumura et al. 2022) is a recent popular method that allows agents to “inherit” other agents’ priorities. Conflict Based Search (Sharon et al. 2015) is another popular method that decoupled the planning problem into two stages. A high-level search resolves collisions between agents by applying constraints while a low-level search finds paths for individual agents that satisfy constraints. There are many extensions to CBS that improve the searches as well as the applied constraints (Barer et al. 2014; Li, Ruml, and Koenig 2021; Li et al. 2021b).
When faced with shorter planning times, methods typically simplify the planning problem to just find partial collision-free paths. Windowed Hierarchical Cooperative A* (Silver 2005) is a windowed variant of Hierarchical Cooperative A* which is essentially a prioritized planner using a backward Dijkstra’s heuristic, and is not complete due to their use of priorities. Rolling Horizon Conflict Resolution (RHCR) applies a rolling horizon for lifelong MAPF planning and replans paths at repeated intervals (Li et al. 2020). RHCR faces deadlock and attempts to combat it by increasing the planning window but still notes that their method is incomplete. Bounded Multi-Agent A* (Sigurdson et al. 2018) proposes that each agent runs its own limited horizon real-time planner considering other agents as dynamic obstacles. However, the method acknowledges that deadlock occurs when agents need to coordinate with one another.
Planning and Improving while Executing (Zhang et al. 2024) is a recent work that attempts to quickly generate an initial full plan using LaCAM (Okumura 2022) and then refines it during execution using LNS2 (Li et al. 2022). However, if a complete plan cannot be found, the method resorts to using the best partial path available, making it incomplete in such situations. The winning submission (Jiang et al. 2024) to the Robot Runners competition, due to the tight planning time constraint, leveraged windowed planning of PIBT with LNS (Li et al. 2021a). They explicitly note deadlock in congestion is a significant challenge.
@@ 2.3 Real-Time Single Agent Search
We leverage ideas from “Real-Time” Search, a single-agent heuristic search problem where, due to limited time constraints, the agent is required to iteratively plan and execute partial paths. Despite repeatedly planning partial paths, Real-Time Search methods are designed to maintain completeness. The main innovation in single-agent real-time search literature is that the agent updates (increases) the heuristic value of encountered states. This prevents deadlock/livelock as states that are repeatedly visited have larger and larger heuristic values which encourages the search to explore other areas. A large variety of real-time algorithms such as LRTA* (Korf 1990), RTAA* (Koenig and Likhachev 2006), and LSS-LRTA* (Koenig and Sun 2009) propose to update the heuristic in different ways.
Section B.1 contains a formal proof of how completeness guarantees hold in Real-Time Search. The core idea is that given a current state $s$, cost function $c$, and a partial path leading to a new state $s^{W}$, we update the heuristic value via a standard bellman update equation $h(s)\leftarrow U(s,s^{W}):=\max(h(s),c(s,s^{W})+h(s^{W}))$. Now if an agent does not reach the goal, it must be stuck in deadlock/livelock and repeatedly visiting some states $s\in\mathcal{S}_{stuck}$. Under certain (achievable) conditions, updating the heuristic can be shown to repeatedly increase $h(s)$ for $s\in\mathcal{S}_{stuck}$ (Korf 1990). At some point then, if the agent is optimally planning to the state $s^{W}$ that minimizes $c(s,s^{W})+h(s^{W})$, it will eventually pick an $s^{W}\notin\mathcal{S}_{stuck}$. Thus most single-agent Real-Time Search algorithms require an optimal planner to maintain completeness.
We note that in the above proof sketch, the planning window $W$ does not matter, nor does how many steps in the partial plan the agent chooses to move (e.g. committing to only one step or executing all $W$ steps). Completeness holds regardless if $W\geq 1$ and the agent moves at least one step.
@@ 3 Windowed-MAPF with Guarantees
This section describes our Windowed Complete MAPF framework, WinC-MAPF, for creating windowed MAPF solvers that guarantee completeness. We leverage two key insights. Our first insight is that we can apply single-agent real-time update ideas to MAPF planning if we interpret the MAPF problem as a single-agent problem in the combined joint configuration space. This allows us to update the heuristics of previously seen states enabling completeness. However, just doing this is ineffective due to the large state space. To this extent, our second insight is that we can leverage MAPF’s agent semi-independence to intelligently update the heuristic value of multiple states, allowing the search to fill in heuristic depressions quickly and exit local minima faster.
@@
Figure 1: The iterative planning, execution, and grouped heuristic updates/penalties for WinC-MAPF.
3.1 Planning in Joint Configuration Space
@@
We can leverage single-agent Real-Time Search literature (Section 2.3) if we view our multi-agent problem in the joint configuration space of all agents.
Here is a quick conceptual but mathematically imprecise summary: We view the $N$ agents in their joint configuration $\mathcal{C}$. At every timestep, we query the action generator to return a valid sequence of actions that leads to a configuration $\mathcal{C}^{\prime}$ which minimizes $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$. We move the agent along the path as well as update the heuristic of $\mathcal{C}$ via $h(\mathcal{C})\leftarrow c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$. We can then prove completeness by directly applying single agent real-time search literature.
@@
We now formally define the above. Given $N$ agents, we define the configuration $\mathcal{C}^{t}=[s_{1}^{t},...s_{N}^{t}]$. At every planning iteration, we query a high-level “action generator” to return a sequence of configurations $\Pi^{0:W}=[\mathcal{C}^{0},...,\mathcal{C}^{W}]$ which minimizes $|\Pi|=|\Pi^{0:W}|+|\Pi^{W:T_{max}}|=\sum_{t=0}^{W-1}c(\mathcal{C}^{t},\mathcal{C}^{t+1})+h_{BD}(\mathcal{C}^{W})$. We define the joint cost and heuristic intuitively, $c(\mathcal{C},\mathcal{C}^{\prime})=\sum_{i=1}^{N}c(s_{i},s_{i}^{\prime})$ and $h_{BD}(\mathcal{C})=\sum_{i=1}^{N}h^{*}_{i}(s_{i})$ (BD = Backward Dijkstra).
Heuristic Penalties Now given the transition (partial path) $\mathcal{C}\rightarrow\mathcal{C}^{W}$ from the AG, we update
$h(\mathcal{C})\leftarrow U(\mathcal{C},\mathcal{C}^{W}):=\max(h(\mathcal{C}),c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W}))$ (1)
@@
$h(\mathcal{C})$ is initially set to $h_{BD}(\mathcal{C})$ and gradually increases as agents visit configurations. We use the term “heuristic penalty” (HP) configuration for visited configurations whose heuristic increased via the update equation to denote how the updated heuristic “penalizes” those configurations and encourages the search to explore others. An HP configuration $\mathcal{C}$ has a nonzero increase from the base heuristic value, i.e. $h(\mathcal{C})=h_{BD}(\mathcal{C})+h_{p}(\mathcal{C})$ where the penalty $h_{p}(\mathcal{C})>0$.
Action Generator Given $h(\mathcal{C})$ which includes states with HPs and without, we want an optimal action generator AG that finds $\operatorname*{arg\,min}_{C{{}^{W}}}c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W})$. If so, we get the same completeness guarantees by directly reusing the single-agent proof. We require an optimal windowed AG for our proof of completeness but conjecture that future work could relax this.
3.2 Disjoint Agent Groups
The framework we have described so far suffers from an obvious issue; the number of configurations $\mathcal{C}$ grows exponentially with the number of agents. Consequently, escaping local minima, which in our context are agents stuck in deadlock/livelock, can require updating the heuristic of an impractical number of states.
Our key idea is therefore to compute and apply heuristic penalties to specific groups of agents instead of on the entire configuration. This significantly speeds up performance as it allows the search to focus on specific “stuck” agents instead of on all agents. Thus given a configuration transition $\mathcal{C}\rightarrow\mathcal{C}^{W}$, we decompose all the agents into disjoint agent groups such that agents between different groups are not interacting with each other. For example, in Figure 1, we would like to determine that the blue and green agents are blocking each other but that the orange agents are not.
Definition 1 (Disjoint Agent Groups).
Given a configuration transition $\mathcal{C}\rightarrow\mathcal{C}^{W}$, and set of disjoint agent groups $\{Gr^{i}\}$, we have the property that for each agent $R^{j}$ with transition $s_{j}\rightarrow s_{j}^{W}$ in disjoint agent group $Gr^{i}$, there cannot exist another agent in a different group $Gr^{k}$ that blocked $R^{j}$ from picking a better path.
Conceptually, this means that each group’s decision to make $\mathcal{C}_{Gr^{i}}\rightarrow\mathcal{C}_{Gr^{i}}^{W}$ is independent of agents in other groups. Instead of computing disjoint groups where agents are not interacting, it is easier to determine coupled agent groups where they could be interacting.
Definition 2 (Coupled Agents).
Given a configuration transition $\mathcal{C}\rightarrow\mathcal{C}^{W}$, an agent $R_{i}$ is coupled with $R_{j}$ if $R_{j}$ prevents $R_{i}$ from choosing a better path or vice-versa.
Note that coupled agents must be in the same disjoint agent group $Gr$. Importantly, disjoint agent groups do not need to solely consistent of coupled agents, i.e. it is okay for extra non-coupled agents (e.g. agents independent of all others) to be in disjoint groups.
From an abstract perspective, we can compute this by iterating through each agent $R_{i}$ which is not on its optimal single-agent path and storing the ids of the other agents which prevented it from picking a better path (there must exist at least one other agent $R_{j}$ otherwise $R_{i}$ could have gone on its optimal path). This builds a dependency graph where agents that share an edge $R_{i}\rightarrow R_{j}$ denote $R_{i}$ blocked by $R_{j}$. We can then find all the disjoint connect components in this dependency graph (e.g. via a DFS) where each disjoint connected component depicts a group of agents $Gr^{i}$ that are blocking each other and are independent from other groups. Instead of updating $h(\mathcal{C})\leftarrow U(\mathcal{C},\mathcal{C}^{W})$, we do $h(\mathcal{C}_{Gr^{i}})\leftarrow U(\mathcal{C}_{Gr^{i}},\mathcal{C}_{Gr^{i}}^{W})$ for each group of agents $Gr^{i}$ at configuration $\mathcal{C}_{Gr^{i}}$.